OpenStudy (goformit100):

A fly lands on one wall of a room. The lower left-hand corner of the wall is selected as the origin of a twodimensional cartesian coordinate system. If the fly is located at the point having coordinates (2.00, 1.00) m, (a) how far is it from the corner of the room? (b) what is its location in polar coordinates?

OpenStudy (vincent-lyon.fr):

Do you have any idea?

OpenStudy (chmvijay):

can u draw the imaginary diagram

OpenStudy (goformit100):

ye i have the idea of 3d here

OpenStudy (goformit100):

Nopope the diagram is in 3d how to draw it in 2d ?

OpenStudy (vincent-lyon.fr):

This is just a 2D problem. You piece of paper is the wall. If you are 2 units to the right and 1 unit up from any origin, the distance is really easy to measure or work out.

OpenStudy (vincent-lyon.fr):

Are you ok now? or do you need more help?

OpenStudy (goformit100):

Ok, I got the answer after your guidance

OpenStudy (goformit100):

Thank you sir

OpenStudy (vincent-lyon.fr):

yw :) do you also have the polar coords?

OpenStudy (goformit100):

No

OpenStudy (vincent-lyon.fr):

Have you drawn this?  θ is not hard to find since you know both sides of the triangle.

OpenStudy (goformit100):

ok

OpenStudy (vincent-lyon.fr):

Remember conversions go: From car to pol: \(r=\sqrt{x^2+y^2}\) \(\theta=\arctan (y/x)\) From pol to car: \(x=r\,\cos\theta\) \(y=r\,\sin\theta\)